Skip to Main Content
PTAGIS logo link back to home page

Interrogation Site Metadata

Search by typing either a site name, site code, organization, or site type.

85M - Eightmile Ck at Fivemile Ck

Status:   Decommissioned
Site Type: Instream Remote Detection System  |  Organization: Oregon Dept. of Fish and Wildlife
General Operational Period: Year Round
Start Date: 3/22/2011 12:00 AM
End Date: 12/3/2018 10:00 AM

INSTREAM DETECTION ARRAY AT THE CONFLUENCE OF EIGHTMILE CREEK AND FIVEMILE CREEK, CONSISTING OF A FS1001M MULTIPLEXING TRANCEIVER, CELL MODEM, AND SIX ANTENNAS. THE ANTENNA CONFIGURATION CONSISTS OF THREE ANTENNAS IN EACH CREEK. EACH ANTENNA SPANS THE CREEK WITHIN THE NORMAL HIGH WATER MARK AND IS SECURED TO THE SUBSTRATE FOR PASS-OVER DETECTION. SITE IS SOLAR-POWERED AND OPERATES YEAR-ROUND. SITE IS ON PRIVATE LAND. Site removed December 2018.

Contacts

Data Steward:

Derrek Faber
Derrek.M.FABER@odfw.oregon.gov
541-296-8026

Location

Coordinates: 45.593295, -121.080635  |  River Km: 309.004.002

Operational History

Operational history describes the period the site was interrogating for PIT tags. The comments provide anecdotal information about the operation of the site during this period. Site operational periods are updated typically at the end of the calendar year.

Year Start Date End Date Comments
2018 1/1/2018 12:00 AM 12/3/2018 10:00 AM
2017 1/1/2017 12:00 AM 1/1/2018 12:00 AM
2016 1/1/2016 12:00 AM 1/1/2017 12:00 AM
2015 1/1/2015 12:00 AM 1/1/2016 12:00 AM
2014 1/1/2014 12:00 AM 1/1/2015 12:00 AM
2013 1/1/2013 12:00 AM 1/1/2014 12:00 AM No interruption in operations
2012 1/1/2012 12:00 AM 1/1/2013 12:00 AM No interruption in operations.
2011 3/22/2011 12:00 AM 1/1/2012 12:00 AM First year of operations.

Configuration

Configuration metadata includes details about the equipment used at the site to detect and record PIT tags. A new configuration sequence record is created when equipment is added or removed from the site and the period of time it was active. A configuration sequence without an End Date is currently active.

Select Configuration Number:
Start Date:  3/20/2011
End Date:  ----

Site Configuration Diagrams:

Equipment History

Select a type to view the history of equipment installed at this site. A missing End Date implies the information about the equipment is current.

Select Equipment Type:
Equipment Code Description Start Date End Date

No Photos

Event Log Summary

This displays the last 25 events, to see the full history, log in to the Dashboard and go to Interrogation Site > Event Logs.

Event Date Event Type Comments
10/13/2016 12:00:00 AM General Antenna # 4, 5 and 6 down today only
10/11/2016 12:00:00 AM General Antenna # 4 down today only, high noise still continues on 5 and 6.
9/19/2016 12:00:00 AM General Antenna # 4 down today only
9/17/2016 12:00:00 AM General Noise reduced on antennas 2 and 3. Antennas 4,5 and 6 up but with high noise.
9/9/2016 12:00:00 AM General Antennas #1, 2, and 3 up with high noise on 2 and 3. Antennas 4-6 still down
7/25/2016 12:00:00 AM SiteDown System down
6/30/2016 12:00:00 AM SiteUp System up
6/29/2016 12:00:00 AM SiteDown System down, high noise on 6/28/2015
1/22/2016 12:00:00 AM SiteUp System up, noise reduced
1/16/2016 12:00:00 AM SiteDown System down
1/8/2016 12:00:00 AM SiteUp System up, with moderate noise
12/28/2015 12:00:00 AM SiteDown System down
12/23/2015 12:00:00 AM SiteUp System up
12/22/2015 12:00:00 AM SiteDown System down
12/13/2015 12:00:00 AM SiteUp System up
12/11/2015 12:00:00 AM SiteDown System down
12/3/2015 12:00:00 AM General System down 12/3/2015 only, high noise on 12/2/2015
11/2/2015 12:00:00 AM General System down 11/2/2015 only, high noise on 11/1/2015
6/29/2015 12:00:00 AM General High noise for one day
3/25/2015 12:00:00 AM SiteUp System up and noise eliminated
3/6/2015 12:00:00 AM SiteDown System down or high noise
2/24/2015 12:00:00 AM SiteUp System up
12/31/2014 1:00:00 AM General Continue to see intermittent power (detection). Ice fog in area, no sun for power.
12/12/2014 1:00:00 AM SiteDown Site Down intermittent power. Ice fog in area, no sun for power.
2/2/2014 1:00:00 AM SiteUp Site Up, power restored.

Archived Event Logs